OpenStudy (anonymous):

need algebra 2 help oh so bad

OpenStudy (anonymous):

What are the roots of 12x = 9 + 5x^2? .... help please will fan and medal

OpenStudy (campbell_st):

rewrite the equation as \[5x^2 - 12x + 9 = 0\] then use the general quadratic formula \[x = \frac{-b \pm \sqrt{b^2 - 4ac}}{2a}\] in your question a =5, b = -12 and c = 9 the roots are complex... hope it helps

OpenStudy (anonymous):

thanks you helped alot